Lilian9667:

I need Help please What is the equation of the line in slope-intercept form? Enter your answer in the boxes. y =__x +__

Timmyspu:

y=mx+b

Timmyspu:

is that what you are looking for

Lilian9667:

I think Numbers go there I think?

Lilian9667:

@ThisGirlPretty I need help or @Hoodmemes

Timmyspu:

I am pretty sure that is your answer

Lilian9667:

hmmm

Lilian9667:

Sorry I forgot the Attachment lol

1 attachment
Timmyspu:

that is what i lerned when i was in 8th grade that y=mx+b

Lilian9667:

Haha This about (Equations from slope Graphs)

Timmyspu:

rise over run

Lilian9667:

That dosen't help dude XD

Timmyspu:

you do the y first and then do the x but if there is not a fraction for the x and only one number the there will always be a 1 even if there is no number at all for the x it will be 1/1

Lilian9667:

ok, @ThisGirlPretty Is he right?

ThisGirlPretty:

i have to refresh my brain with this~ It's been so long since I learned how to do this

Lilian9667:

Oh ok lol

ThisGirlPretty:

Do you understand how to do the problem?

Timmyspu:

ya i do and that is the correct awnser

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ThisGirlPretty Do you understand how to do the problem? \(\color{#0cbb34}{\text{End of Quote}}\) Me?

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu ya i do and that is the correct awnser \(\color{#0cbb34}{\text{End of Quote}}\) It is?

Lilian9667:

If you were talking to me I don't even know lol

Timmyspu:

ya i was

XioGonz:

y=mx+b is correct \[\frac{ Rise }{ Run }\]

Lilian9667:

Hey! no other subjects

Timmyspu:

i smary

Timmyspu:

smart*

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz y=mx+b is correct \[\frac{ Rise }{ Run }\] \(\color{#0cbb34}{\text{End of Quote}}\) Ohh ok That's what Timmyspu was talking about lol

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz y=mx+b is correct \[\frac{ Rise }{ Run }\] \(\color{#0cbb34}{\text{End of Quote}}\) Ohh ok That's what Timmyspu was talking about lol \(\color{#0cbb34}{\text{End of Quote}}\) Yes

Lilian9667:

I mean that is the Answer.

XioGonz:

Yes unless you have other numbers that you need to plug into the equation.

Lilian9667:

No no I don't

XioGonz:

Ah, okay then yes that's the answer.

Lilian9667:

Wow lol

1 attachment
Lilian9667:

Thanks you for the 60% XD

dude:

Yeah so you the m is the rise/run The b is the value in the y axis

1 attachment
XioGonz:

Hey I asked if you had other numbers you need to plug in

dude:

4 is in the y axis in this case

XioGonz:

There was a graph?

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Hey I asked if you had other numbers you need to plug in \(\color{#0cbb34}{\text{End of Quote}}\) Well i didn't know what ya think I'm the one who's asking the Question around here lol

Lilian9667:

Yeah ya didn't see it

XioGonz:

Nope, I guess it's my fault then, I apologize.

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude Yeah so you the m is the rise/run The b is the value in the y axis \(\color{#0cbb34}{\text{End of Quote}}\) Thank you

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Nope, I guess it's my fault then, I apologize. \(\color{#0cbb34}{\text{End of Quote}}\) It's ok people makes mistakes

Timmyspu:

ya lol

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Nope, I guess it's my fault then, I apologize. \(\color{#0cbb34}{\text{End of Quote}}\) It's ok people makes mistakes \(\color{#0cbb34}{\text{End of Quote}}\) I'll be more careful next time

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Nope, I guess it's my fault then, I apologize. \(\color{#0cbb34}{\text{End of Quote}}\) It's ok people makes mistakes \(\color{#0cbb34}{\text{End of Quote}}\) I'll be more careful next time \(\color{#0cbb34}{\text{End of Quote}}\) Ok Imao

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu smart* \(\color{#0cbb34}{\text{End of Quote}}\) And you said you were smart๐Ÿ˜‚๐Ÿ˜‚